OpenStudy (anonymous):

10) At takeoff, a plane flies at an angle of 10° with the runway. After it has traveled a ground distance of 2,800 feet, find the vertical distance the plane has gained from takeoff. Round your answer to the nearest foot. A) 486 ft B) 494 ft C) 2,757 ft D) 15,880 ft

jimthompson5910 (jim_thompson5910):

tan(10) = x/2800 2800*tan(10) = x 2800*0.1763 = x 493.64 = x x = 493.64 Rounding to the nearest foot, the answer is 494 feet

OpenStudy (anonymous):

this is another trig application problem. construct a triangle by: the answer can be found by tan(10) = opp / adj = height / 2800. so opp = tan(10) * 2800 = about 494